Back to the NIKA2 main page

Daily Reports

Feb 1st, Tuesday

16:00-21:00; Ioannis: During maintenance, SN worked with MC on the problem of Ar1 resonances that show peaks around 3 Hz, 6 Hz and 9 Hz. They identified the Ar1 synthesizer as the source of the problem (by swapping the Ar1 and Ar3 synthetizers) and then replaced the faulty synthesizer with the last spare one. The problem is not visible anymore in the traces. We start the NIKA2 tuning after maintenance at 16:30 but we have a problem with Ar3 resonances. We restart several times the acquisition but we cannot tune Ar3 most of the times. After investigating numerous factors with JLS and contacting SN, we suspect that the increased temperature in the receiver cabin (up to 24 ºC) may be causing these problems. We turn off the Ar3 synthesizer, restart the temperature control in the cabin and use a fan for additional cooling of the crates (mainly 3). Around 20:30 the temperature in the cabin drops around 20 ºC. We restart the acquisition and Ar3 behaves normally (bad kids: 43/7/48 for Ar1/Ar2/Ar3), so we start with the observations. After pointing on Uranus we stop at 21:15 due to high wind. Observations restart at 22:00. Pointing and focus on Uranus but the beam is a bit elongated (17"x20" at 2mm and 11"x14" at 1mm), most likely because the antenna temperature evolution shows that the cooling of the structure is slower than usual today (highest temperature peak in the last few weeks). Will have to wait a bit longer for the antenna to fully relax.

21:00-03:00; Guilaine: Exceptional tau (0.15) but very bad shift. Electronic problems at the beginning. Then windy. Start observing Uranus at 22:10. Series of pointing/focus. Beams way too bad (today was the hottest day since 1 month). Change source to the very bright 0316+413 (since Uranus at too low elevation). New series of pointing/focus. Skydip [20220201s158]. 23:28, beams of 12.13 – 17.92 – 12.29. Launch 1 repet on source NGC1333-C7 (18 minutes), which is very close to 0316+413. Then again series of pointing/focus. Although the focus was at expected value, very degraded beams. 1:06: do a "retune". And restarting the whole acquisition (too much bad KIDS in array1). 2:00: pointing scan on 0552+398. Beams of 12.78, 18.33, 12.90. Got to project 137-21 to observe a low priority target (WB89-640). Scan 14 aborted at 03:12 since array 3 was lost (lost before 03:05, I don’t know exactly when, cannot see the acquisition screen remotely; scan 12 reduced with nk_rta and doesn’t seem weird).

Feb 2nd, Wednesday

03:00-09:00; Juan: At the beginning of the shift array 3 showed problems again -operator went up to the cabin to restart and cool down the synthetizer. We finally manage to observe for 2 hours or so, 2.5 repetions on PSZ2G122

09:00-15:00; Ioannis: At 09:30 SN arrived at the telescope and we tried to resolve the issue with the Ar3 synthesizer. First we replaced it with the previous one but unfortunately didn't work. Then we tried the synthesizer used for VLBI and fortunately worked and Ar3 was stable again. PIs are noted that NIKA2 observations between scans 20220101s128 and 20220102s24 may be affected by the unstable behavior of Ar3 synthesizer and are advised to proceed with the data reduction with caution. After the replacement of the Ar3 synthesizer, we restart with NIKA2 observations around 11:00 (good tuning with 38/6/47 bad kids for Ar1/Ar2/A3), but unfortunately the beam is relatively large and elongated even after several focus attempts. We take this opportunity and do some polarization tests with NP, MC, JMP, AB, PA, DJ and AS, in view of the upcoming B-FUN observations this Friday. We made sure that HWP operated as expected, go through the steps required to prepare the acquisition software for polarimetric observations and make sure that the IDL pipeline (called by "idl_polar" for NIKA2-Pol) is able to reduce pointings, focus, skydips and science observations. IMBFITS are created for polarization data but the PIIC monitors are not analyzing them at the moment. The tests last until the weekly NIKA2 meeting at 15:00.

15:00-21:00; Isabelle & Aurore: [IM: We restart with NIKA2 observations at 16:00. However, the beam is still too large and elongated to start with the science observations (12"x20" for 1mm and 17"x22" for 2mm). The situation improves slightly around 17:00 and we do a calib_1scan on Uranus [20220102s113] and then a skydip [20220102s114]. At 17:30 the wind gets too high to continue (> 20 m/s) and we stop the antenna.] Stopped throughout the shift due to high wind.

21:00-03:00; Matthieu: The opacity was fantastic (tau~0.1) during the 9pm-3am shift. However, the wind (15-25 kms) decided otherwise.

Feb 3rd, Thursday

03:00-09:00; Corentin & Miren: Stopped throughout the shift due to high wind.

09:00-15:00; Ioannis: After a long pause due to high wind, we restart NIKA2 observations around 12:00. Good tuning with 34/5/45 bad kids for Ar1/Ar2/Ar3. Pointing and focus on 1633+382 with small and symmetrical beams in both bands. Then observations of L694 (2 repetitions) for project 109-21. Then we move to project 137-21 and observed 19383+2711 (second attempt because first time it was interrupted).

15:00-21:00; Isabelle & Aurore: Good weather conditions and stability, slightly distorted beams at the beginning, improving. Eight scans on L1521E from 15:30 UT to 18:50 UT. calib1scan [20220203s44] and skydip [20220203s45] performed at the beginning of the shift (15:20 UT). [IM: At 19:50 we performed EMIR observations for about 1.5 hrs, as requested by RN and CK].

21:00-03:00; Ioannis: Due to degrading weather, we continued with EMIR and observed for the backup project 170-21.

Feb 4th, Friday

03:00-09:00; Laurence: No NIKA2 observations due to poor weather; tau225 of about 0.6 (with variations from 0.55 to 1). [IM: EMIR observations until 08:00. Then stopped because the weather was too bad even for that.]

09:00-15:00; Ioannis: Bad and unstable weather (tau225 between 0.5 and 1). No observations.

15:00-21:00; Ioannis: Bad weather continues. No observations.

21:00-03:00; Ioannis: Bad weather continues plus some rain and snow. No observations.

Feb 5th, Saturday

03:00-09:00; Ioannis: Bad weather continues plus some rain and snow. No observations.

09:00-15:00; Ioannis: Fog and some snow throughout the shift No observations.

15:00-21:00; Ioannis: Fog and some snow throughout the shift No observations.

21:00-03:00; Ioannis: Opacity stabilized around 0.4-0.6. EMIR backup observations of project 170-21 since 22:00.

Feb 6th, Sunday

03:00-09:00; Miguel & Ioannis: Tau225 oscillating between 0.4 and 0.6. EMIR backup observations (170-21) throughout the night.

09:00-15:00; Miguel & Ioannis: Weather deteriorated around 12:00 and we stop EMIR backup observations.

15:00-21:00; Jean-François & Ioannis: Tried to restart observations around 17:30, initially with EMIR. The local weather conditions are still quite unstable for that. The telescope is continuously inside and outside of a cloud. After the weather stabilized, we start with NIKA2 observations around 19:30. Good tuning with 30/4/43 bad kids for Ar1/Ar2/Ar3. However, the beam is still a bit elongated (~17"x20" at 2mm) and we get large pointing offsets of about +15" and -15" for azimuth and elevation, respectively. We checked different possible sources for this effect, including the wobbler system at M2 and the optical path to NIKA2, but the large offsets persist. We will continue the investigation for this effect.

21:00-03:00; Philippe, Nico & Hamza: The weather conditions have improved significantly with stable opacity around 0.15. We change to the NIKA2-pol mode (HWP setup, DAQ restart with run70_pol.ini file and retuning). Pointing and focus on Uranus. The beam elongation gradually improves but the pointing offset remain large (~15" and -20"). Nevertheless, the pointing variability is very small (<3") and we start with the B-FUN observations, focusing on OMC-1 and using Uranus for calibrations. We did 2 calib_1scan on Uranus [20220206s178, 20220206s190], observed OMC-1 and NGC2264 (B-FUN targets) as well as pointing sources (quasars) with calib_1scan [20220206s192, 20220206s194] to try the CLUD polarization calibration method. The pointing offsets continue to be large and show also large variations between pointing scans on different sources. Focus was stable and opacity low ~0.1.

Feb 7th, Monday

03:00-09:00; Emmanuel and Aishwarya: Opacity was good and very stable at ~<0.11 during the whole shift. However, due to the telescope pointing problem, we were limited to observe between elevation range of 45+-10 and the beam was wide throughout the shift (~>18" @2mm and ~11" @ 1mm) with an elongated shape. We performed one repetition on PSZ2G081 (199-16). After making a pointing on the bright source 1253-055 for the beam map we proceeded to project 199-16 since the pointing was not too good for the beam map. We proceeded to make three reps on PSZ2G133 (199-16). Pointing was performed and the opacity was monitored between each rep. Towards the end of the shift, we performed a final pointing which again, was not good for a calibration scan and ended with a sky dip [20220207s64].

09:00-15:00; Ioannis: We investigated further the large pointing offset issue at the beginning of the shift. After checking different factors, we decided to change back to EMIR setup to see if this is a common problem also for EMIR. After changing to EMIR setup and then back to the NIKA2 setup in the receiver cabin, the pointing offsets are not so large anymore. For an elevation range between 20º and 64º, we get Δaz = -4.9” +/- 1.1” and Δel = -2.5” +/- 1.8” for EMIR and Δaz = -3.4” +/- 1.6” and Δel = -5.6” +/- 1” for NIKA2. It is not clear what caused the issue, maybe some problem during the NIKA2 mirrors deployment yesterday, but we are not sure yet. Unstable and elongated beam and variable focus until the end of the shift, most likely due to (strong) temperature gradients across the antenna from the direct sunlight.

15:00-21:00; Aurore: Good weather conditions and stability, focus values were rather stable, but beams distorted at the beginning (Uranus looked like Mickey Mouse, or was very elongated, sometimes 16"x30" at 1 mm!), improving during the shift. Eight scans on L1521E from 15:30 UT to 19:15 UT. calib_1scan [20220207s140] and skydip [20220207s141] performed at the beginning of the shift, however beam was not great at the time. IM: I took over around 20:15 for some calibrations. The beam size and shape are restored to their nominal values (17.3"x18.1" at 2mm, 10.1"x12.1" at 1mm) and atmosphere is stable (baseline variations: +/- 2 Jy/beam at 2mm, +/- 5 Jy/beam at 1mm). I do a calib_1scan on Uranus [20220207s191], a skydip [20220207s192] and a beam map on Uranus [20220207s194].

21:00-04:00; Philippe & Hamza: Very good weather conditions with tau225 ~ 0.18. Pointing much improved compared to the previous night. We are started with NIKA2-Pol around 21:30, following an (unpolarized) beam map on Uranus [20220207s194]. With NIKA2-Pol, we took a calib_1scan on Uranus [20220207s201], 1 sequence (of 3 maps) on OMC1, 2 sequences on OMC2, and 3 sequences on NGC2264 for the B-FUN project. We also observed a polarized skydip [20220208s19]. At the end of the shift, after 03:00, we took 1 calib_1scan on 3C273 [20220208s26] and 1 calib_1scan on 3C279 [20220208s27] for comparaison with X-Pol (starting at 04:00).

Feb 8th, Tuesday

04:00-10:30; Ioannis: EMIR polarimetric observations for POLAMI LP (087-21).

DailyReportsNika2Pool010222 (last edited 2022-02-08 04:54:12 by NikaBolometer)